# Quantum Cosmology, Decoherence, and the Emergence of Classical Spacetime

### v1

@@ Abstract
We analyze the emergence of classical cosmological spacetimes in quantum cosmology by computing the reduced density matrix for long-wavelength curvature perturbations. Starting from standard Hartle–Hawking and tunneling boundary conditions, we emphasize that semiclassical WKB structure and inflationary squeezing do not by themselves yield classicality. Tracing over unobserved degrees of freedom and using the influence functional formalism, we derive the decoherence functional for superhorizon curvature modes during inflation. For a light massive environmental scalar field in the Bunch–Davies vacuum, we obtain an explicit noise kernel and show how a nonzero mass regulates the infrared behavior. We then evaluate decoherence under horizon-based and EFT-motivated coarse grainings, finding efficient suppression of interference between macroscopically distinct perturbation histories in both cases. The analysis clarifies the distinct roles of boundary conditions (branch amplitudes) and decoherence (classical branch selection) and yields an emergent cosmological arrow of time through environment-induced entanglement.
Quantum cosmology, Decoherence, Influence functional, Inflation, Emergence of classicality
@@ I Quantum Cosmology Before Classicality
I.1 Hilbert’s Sixth Problem and Emergent Classicality
Hilbert’s Sixth Problem calls for the axiomatization of physical theories [1] and, in particular, for a precise understanding of how macroscopic classical laws emerge from underlying microscopic dynamics. In modern language, this problem encompasses the emergence of irreversible, probabilistic, and effectively classical behavior from fundamentally quantum and reversible laws. While the problem was originally posed in the context of mechanics and probability theory, its scope naturally extends to quantum field theory, statistical mechanics, and cosmology.
In quantum cosmology, Hilbert’s Sixth Problem takes on a sharpened form: how do classical spacetime, classical matter fields, and classical stochastic perturbations arise from a quantum theory of gravity and quantum fields in curved spacetime? The Wheeler–DeWitt equation and its associated boundary conditions define a universal quantum state, but they do not by themselves explain why observations are accurately described by classical cosmology.
Recent formulations of Hilbert’s Sixth Problem emphasize the role of coarse graining, statistical limits, and emergent descriptions in bridging microscopic and macroscopic theories. In this spirit, the influence functional and decoherence framework provides a concrete and calculable realization of Hilbert’s program in cosmology. By explicitly tracing over unobserved degrees of freedom, one obtains effective dynamics for relevant variables that are irreversible, stochastic, and classical in precisely the sense required by Hilbert’s Sixth Problem.
The approach adopted in this work is directly motivated by these ideas and by recent analyses emphasizing Hilbert’s Sixth Problem as a guiding principle for foundational questions in physics [13]. While Ref. [13] concerns classical limits in kinetic theory, we cite it here as a representative modern implementation of Hilbert VI reasoning—deriving irreversible macroscopic equations from reversible microdynamics—and we apply the same logic to the quantum-to-classical transition in cosmology via coarse graining and decoherence. In particular, we build on the viewpoint that decoherence and coarse graining are not interpretational add-ons, but essential structural elements required to derive classical laws from quantum dynamics. The results presented below implement this program explicitly for cosmological perturbations during inflation, thereby providing a concrete cosmological realization of Hilbert’s Sixth Problem.
@@ I.1 Hilbert’s Sixth Problem and Emergent Classicality
defined over three-geometries and matter field configurations. Two canonical proposals for specifying this wavefunction are the no-boundary proposal of Hartle and Hawking and the tunneling proposal of Vilenkin [2, 3]. In practice, both are implemented in a truncated configuration space consisting of minisuperspace variables, such as the scale factor $a$, together with quantum perturbations.
@@
where $\alpha$ labels distinct semiclassical branches. While this structure allows approximate classical trajectories for $a(t)$ and approximate Schrödinger evolution for perturbations along each branch, the total state remains a coherent superposition of macroscopically distinct geometries. WKB behavior alone therefore does not explain the emergence of classical spacetime, since the universal wavefunctional in Eq. (1) remains a coherent superposition of macroscopically distinct geometries.
@@
Observations necessarily access only a subset of degrees of freedom. The physically relevant object is thus the reduced density matrix obtained by tracing over unobserved variables, defined explicitly in Eq. (3). Classical spacetime can emerge only if this reduced density matrix becomes approximately diagonal in an appropriate basis [4, 5, 8].
@@ II Reduced Density Matrix and Influence Functional
Figure 1: Schematic roadmap (adapted from standard decoherence frameworks, e.g. influence-functional constructions of the Feynman–Vernon type). Quantum-cosmological boundary conditions define the universal state $\Psi$. Tracing over unobserved degrees of freedom yields a reduced density matrix for long-wavelength perturbations. The influence functional encodes dissipation and noise; the associated decoherence functional suppresses interference between macroscopically distinct histories. Classical stochastic descriptions and an operational arrow of time emerge after coarse graining, with the detailed rate depending only weakly on the coarse-graining prescription.
@@
The time evolution of $\rho_{L}$ can be written using the Schwinger–Keldysh formalism [6, 7]. Integrating out the environment defines an influence functional, whose general structure is encoded in the influence action Eq. (4) $\mathcal{F}=\mathrm{e}^{\mathrm{i}S_{\mathrm{IF}}}$, where the influence action admits the general form
$S_{\mathrm{IF}}=\int\mathrm{d}^{4}x\,\mathrm{d}^{4}x^{\prime}\left[\zeta_{\Delta}(x)D(x,x^{\prime})\zeta_{c}(x^{\prime})+\frac{\mathrm{i}}{2}\,\zeta_{\Delta}(x)N(x,x^{\prime})\zeta_{\Delta}(x^{\prime})\right]+\cdots.$ (4)
Here $D$ is a real, causal dissipation kernel and $N$ is a real, positive semidefinite noise kernel. Decoherence is governed by the imaginary term involving $N$.
The suppression of off-diagonal elements of the reduced density matrix is controlled by the decoherence functional defined in Eq. (5)
@@
When $\Gamma\gg 1$ for macroscopically distinct configurations (Eq. (5)), the reduced density matrix becomes approximately diagonal.
@@ III Explicit Noise Kernel for a Light Environment Field
We consider an environment consisting of a light scalar field $\sigma$ with mass $m_{\sigma}\lesssim H$ in the Bunch–Davies (BD) vacuum [9, 8]. The leading effective interaction with the long-wavelength curvature perturbation is modeled as
$S_{\mathrm{int}}=g\int\mathrm{d}\eta\,\mathrm{d}^{3}x\;a^{4}(\eta)\,\zeta_{L}(\eta,\mathbf{x})\,\sigma^{2}(\eta,\mathbf{x}),$ (6)
with $a(\eta)=-1/(H\eta)$. Such a coupling captures the universal fact that long-wavelength curvature perturbations modulate local volumes and hence the dynamics of additional sectors.
For a Gaussian environment state, the noise kernel appearing in Eq. (5) entering the influence action is [6, 7]
$N(x,x^{\prime})=g^{2}\,\mathrm{Re}\left[G^{>}_{\sigma}(x,x^{\prime})^{2}\right],$ (7)
@@
$u_{k}(\eta)=\frac{\sqrt{\pi}}{2}e^{\mathrm{i}(\nu+1/2)\pi/2}H(-\eta)^{3/2}H^{(1)}_{\nu}(-k\eta),\qquad\nu=\sqrt{\frac{9}{4}-\frac{m_{\sigma}^{2}}{H^{2}}},$ (9)
@@
Substituting Eqs. (8)–(9) into Eq. (7) yields the explicit convolution
$N(\eta,\eta^{\prime};\mathbf{k})=g^{2}\,\mathrm{Re}\left[\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}u_{q}(\eta)u^{*}_{q}(\eta^{\prime})\,u_{|\mathbf{k}-\mathbf{q}|}(\eta)u^{*}_{|\mathbf{k}-\mathbf{q}|}(\eta^{\prime})\right].$ (11)
@@
$N(\eta,\eta^{\prime};k_{L})\simeq g^{2}\,\mathrm{Re}\left[\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\left(u_{q}(\eta)u^{*}_{q}(\eta^{\prime})\right)^{2}\right].$ (12)
@@
$N(\eta,\eta^{\prime};k_{L})\propto g^{2}H^{4}(-\eta)^{2\Delta}(-\eta^{\prime})^{2\Delta}\int\mathrm{d}q\,q^{-4+4\Delta}.$ (13)
Thus a nonzero mass $m_{\sigma}\neq 0$ softens the infrared behavior and renders the accumulation of decoherence physically controlled by the finite duration and size of the inflationary patch, rather than by a formal divergence. This explicit kernel provides the basis for the coarse-grained estimates discussed in the next section.
IV Coarse Graining and Emergent Classical Dynamics

### v2

# The Cosmological Arrow of Time from Inflationary Branch Decoherence

@@ Abstract
We analyze the emergence of classical cosmological spacetimes in quantum cosmology by computing the reduced density matrix for long-wavelength curvature perturbations. Starting from standard Hartle–Hawking and tunneling boundary conditions, we emphasize that semiclassical WKB structure and inflationary squeezing do not by themselves yield classicality. Tracing over unobserved degrees of freedom and using the influence functional formalism, we derive the decoherence functional for superhorizon curvature modes during inflation. For a light massive environmental scalar field in the Bunch–Davies vacuum, we derive the convolution structure and superhorizon scaling of the noise kernel and show how a nonzero mass softens the infrared behavior. We then evaluate decoherence under horizon-based and EFT-motivated coarse grainings, finding efficient suppression of interference between macroscopically distinct perturbation histories in both cases. We further derive the branch-overlap factor $|\mathcal{D}_{k}(z)|$ explicitly from the Bunch–Davies mode functions in the expanding-versus-contracting branch formalism, obtaining the exact closed form $|\mathcal{D}_{k}(z)|=[z^{2}/(z^{2}+1)]^{1/4}$ (massless limit) and the superhorizon power law $|\mathcal{D}_{k}(z)|\sim z^{\nu}$ (massive fields). Numerical evaluation shows that the geometric (coupling-independent) branch-decoherence functional crosses unity within $\approx 0.5$ e-folds, growing irreversibly thereafter, providing a derived account of why expanding cosmological histories acquire robust classical records while contracting ones do not. The analysis clarifies the distinct roles of boundary-condition amplitudes and environment-induced decoherence and connects the emergent cosmological arrow of time directly to inflationary squeezing.
pacs:
98.80.Qc, 03.65.Yz, 04.60.Kz, 98.80.Cq
@@ I Quantum cosmology before classicality
I.1 Decoherence and the classical limit in quantum cosmology
Quantum cosmology seeks to explain how a universal quantum state gives rise to the effectively classical spacetime described by cosmological observations. The Wheeler–DeWitt equation and its associated boundary conditions define a wavefunctional of geometry and matter, but they do not by themselves explain why observations are well described by classical backgrounds with stochastic perturbations.
This issue is sharpened during inflation. Semiclassical WKB structure can identify approximate background trajectories, and squeezing can make correlation functions appear classical, yet neither effect by itself suppresses phase coherence between macroscopically distinct alternatives. The quantum-to-classical transition therefore requires a dynamical mechanism acting on the reduced state of the observable degrees of freedom.
The natural framework for that problem is coarse graining and environment-induced decoherence. By tracing over inaccessible degrees of freedom one obtains an influence functional for the long-wavelength variables, with dissipation and noise kernels that govern the loss of phase coherence and the onset of effectively classical stochastic dynamics. In this sense, classical spacetime is not put in by hand but emerges from a reduced description of a larger quantum state.
The aim of this paper is to make the dynamical origin of classical spacetime explicit by deriving the decoherence functional for inflationary curvature perturbations from the underlying action. Our focus is not only on the general claim that decoherence occurs, but on three more specific questions: how the effective long-mode coupling to a spectator sector arises, how the decoherence rate depends on the coarse graining, and how branch decoherence clarifies the relation between quantum-cosmological boundary conditions and classical spacetime histories.
@@ I.1 Decoherence and the classical limit in quantum cosmology
defined over three-geometries and matter field configurations. Two canonical proposals for specifying this wavefunction are the no-boundary proposal of Hartle and Hawking and the tunneling proposal of Vilenkin [1, 2]. The semiclassical branch structure of these wavefunctions and their relation to classical histories has been studied within the decoherent-histories framework [9, 8]. In practice, both proposals are implemented in a truncated configuration space consisting of minisuperspace variables, such as the scale factor $a$, together with quantum perturbations.
@@
where $\alpha$ labels distinct semiclassical branches. While this structure allows approximate classical trajectories for $a(t)$ and approximate Schrödinger evolution for perturbations along each branch, the total state remains a coherent superposition of macroscopically distinct geometries. WKB behavior alone therefore does not explain the emergence of classical spacetime, since the universal wavefunctional (1) remains a coherent superposition of macroscopically distinct geometries.
@@
Observations necessarily access only a subset of degrees of freedom. The physically relevant object is thus the reduced density matrix obtained by tracing over unobserved variables, defined explicitly in equation (3). Classical spacetime can emerge only if this reduced density matrix becomes approximately diagonal in an appropriate basis [3, 4, 7, 16].
Recent work has further emphasized the role of environment-induced decoherence in inflationary cosmology. In particular, Burgess, Holman, and Hoover showed that interactions between long-wavelength perturbations and subhorizon environmental degrees of freedom can generate efficient decoherence of inflationary fluctuations, even when the underlying dynamics remains unitary [12]. Their analysis demonstrated that the decoherence rate is controlled by the coupling between system and environment as well as the rapid growth of physical phase space during inflation. Our treatment is complementary to these approaches: rather than focusing primarily on phenomenological estimates within an inflationary EFT, we derive the effective spectator-field coupling from the underlying action, construct the associated influence-functional noise kernel, and use the resulting branch-overlap structure to separate the role of boundary-condition amplitudes from the dynamical emergence of classical histories.
What is new here.— While decoherence of inflationary perturbations has been discussed in earlier work (e.g. [12, 13]), the present analysis makes five specific contributions. (i) We derive the effective long-wavelength spectator coupling directly from the underlying covariant action in comoving gauge (rather than postulating it within an EFT) and identify the corresponding influence-functional noise-kernel structure. (ii) We compare horizon-based and effective-field-theory coarse grainings within a common framework and show that both lead to robust growth of the decoherence functional. (iii) We evaluate the branch-overlap factor $|\mathcal{D}_{k}(z)|$ explicitly using the Bunch–Davies mode functions, obtaining the exact closed form $|\mathcal{D}_{k}(z)|=[z^{2}/(z^{2}+1)]^{1/4}$ in the massless limit and the power law $|\mathcal{D}_{k}(z)|\propto z^{\nu}$ for massive fields. To our knowledge this explicit evaluation has not appeared previously: earlier works characterised classicality through squeezing parameters (Polarski and Starobinsky [10]), reduced-density-matrix elements (Kiefer and Polarski [13]), or noise-kernel rates (Burgess, Holman and Hoover [12]), but did not extract the closed-form branch-overlap factor. (iv) We derive the dissipation kernel explicitly and show that it is consistent with the de Sitter fluctuation-dissipation relation at Gibbons–Hawking temperature $T_{\mathrm{GH}}=H/(2\pi)$; the resulting Langevin equation recovers Starobinsky’s stochastic inflation in the superhorizon limit, in agreement with the standard derivation. (v) We explicitly separate boundary-condition amplitudes from dynamical branch-overlap suppression, thereby clarifying the different roles of the Hartle–Hawking and Vilenkin proposals versus environment-induced decoherence in generating the cosmological arrow of time.
Figure 1: Schematic roadmap of the emergence of classical cosmological histories from quantum-cosmological boundary conditions via influence-functional decoherence [5, 6]. Boundary conditions define the universal state $\Psi$. Tracing over unobserved degrees of freedom yields a reduced density matrix for long-wavelength perturbations. The influence functional encodes dissipation and noise; the associated decoherence functional suppresses interference between macroscopically distinct histories. Classical stochastic descriptions and an operational arrow of time emerge after coarse graining, with the detailed rate depending only weakly on the coarse-graining prescription.
@@ II Reduced density matrix and influence functional
The time evolution of $\rho_{L}$ can be written using the Schwinger–Keldysh formalism [5, 6, 21, 22]. Integrating out the environment defines an influence functional whose general structure is encoded in the influence action $\mathcal{F}=\mathrm{e}^{\mathrm{i}S_{\mathrm{IF}}}$, where
$S_{\mathrm{IF}}=\int\mathrm{d}^{4}x\,\mathrm{d}^{4}x^{\prime}\Bigl[\zeta_{\Delta}(x)\,D(x,x^{\prime})\,\zeta_{c}(x^{\prime})+\frac{\mathrm{i}}{2}\,\zeta_{\Delta}(x)\,N(x,x^{\prime})\,\zeta_{\Delta}(x^{\prime})\Bigr]+\cdots.$ (4)
Here $D$ is a real, causal dissipation kernel and $N$ is a real, positive-semidefinite noise kernel. Decoherence is governed by the imaginary term involving $N$.
The suppression of off-diagonal elements of the reduced density matrix is controlled by the decoherence functional
@@
When $\Gamma\gg 1$ for macroscopically distinct configurations, the reduced density matrix becomes approximately diagonal.
A note on two distinct decoherence functionals. This paper computes $\Gamma$ in two complementary ways that yield different e-fold scales and should not be confused. In section IV we evaluate the geometric branch-decoherence functional $\Gamma_{+-}=-\sum_{k}\ln|\mathcal{D}_{k}|$, which measures how differently the expanding and contracting semiclassical backgrounds squeeze each environmental mode; this quantity is independent of the spectator coupling $\hat{\lambda}$ and crosses unity within $\approx 0.5$ e-folds. In sections III and III.2 we evaluate the interaction-dependent noise-kernel functional, which measures the rate at which a specific spectator sector suppresses off-diagonal elements of the reduced density matrix for the system perturbation $\zeta_{L}$; this crosses unity at several to $\sim 10$ e-folds depending on $\hat{\lambda}$. The two quantities are not competing predictions for the same process: the first isolates the geometric asymmetry between cosmological branches, while the second quantifies how system–environment interactions exploit that asymmetry. Both confirm irreversible classicalisation well within inflation.
@@ III Explicit noise kernel for a light environment field
We consider an environment consisting of a light spectator scalar field $\sigma$ with mass $m_{\sigma}\lesssim H$ in the Bunch–Davies (BD) vacuum [10, 7, 14, 15]. Rather than postulating the system–environment coupling, we derive its leading long-wavelength form from the covariant spectator action
$S_{\sigma}=-\frac{1}{2}\int\mathrm{d}^{4}x\,\sqrt{-g}\left[g^{\mu\nu}\partial_{\mu}\sigma\,\partial_{\nu}\sigma+m_{\sigma}^{2}\,\sigma^{2}\right].$ (6)
Working in comoving gauge, where the inflaton fluctuation is set to zero and the scalar metric perturbation is carried by $\zeta$,
$h_{ij}$ $=a^{2}e^{2\zeta}\delta_{ij},$
$N$ $=1+\frac{\dot{\zeta}}{H}+\mathcal{O}(\zeta^{2},\partial^{2}),\qquad N_{i}=\partial_{i}\psi,$ (7)
the spectator action becomes, using the ADM decomposition,
$S_{\sigma}=\frac{1}{2}\int\mathrm{d}t\,\mathrm{d}^{3}x\,N\sqrt{h}\Bigl[N^{-2}\!\left(\dot{\sigma}-N^{i}\partial_{i}\sigma\right)^{\!2}-h^{ij}\partial_{i}\sigma\,\partial_{j}\sigma-m_{\sigma}^{2}\,\sigma^{2}\Bigr].$ (8)
Expanding to first order in the long-wavelength curvature perturbation gives
$S_{\sigma}=S_{\sigma}^{(0)}+S_{\sigma\zeta}^{(3)}+\cdots,$ (9)
with
$S_{\sigma\zeta}^{(3)}=\frac{1}{2}\int\mathrm{d}t\,\mathrm{d}^{3}x\,a^{3}\left[\left(3\zeta+\frac{\dot{\zeta}}{H}\right)\dot{\sigma}^{2}-\left(\zeta+\frac{\dot{\zeta}}{H}\right)\frac{(\partial_{i}\sigma)^{2}}{a^{2}}-\left(3\zeta+\frac{\dot{\zeta}}{H}\right)m_{\sigma}^{2}\,\sigma^{2}\right]+\cdots,$ (10)
where the ellipsis denotes terms proportional to $N_{i}$ and slow-roll-suppressed constraint corrections. As discussed in Maldacena’s detailed treatment of the comoving-gauge constraints [14], the shift $N_{i}$ is determined by the momentum constraint and its contribution at this order enters only through constraint-suppressed terms, which is why the $N_{i}$ sector can be consistently grouped with the slow-roll-suppressed corrections in equation (10). Equation (10) makes clear that the coupling is not an arbitrary ansatz: a long-wavelength $\zeta_{L}$ locally rescales the spectator Hamiltonian density. For the superhorizon system mode of interest, $\dot{\zeta}_{L}/H=\mathcal{O}(\epsilon)\zeta_{L}$ is slow-roll suppressed, and if the spectator is light and dominated by its mass term on sufficiently long wavelengths, equation (10) reduces to the local form
$S_{\mathrm{int}}^{\mathrm{eff}}=-\frac{3}{2}\,m_{\sigma}^{2}\int\mathrm{d}t\,\mathrm{d}^{3}x\,a^{3}\,\zeta_{L}\,\sigma^{2}\equiv-\int\mathrm{d}t\,\mathrm{d}^{3}x\,a^{3}\,\lambda_{\mathrm{eff}}\,\zeta_{L}\,\sigma^{2},$ (11)
with $\lambda_{\mathrm{eff}}=\tfrac{3}{2}m_{\sigma}^{2}$ in the minimal spectator model and with additional $\mathcal{O}(\epsilon H^{2})$ or derivative corrections in more general EFTs of inflation. In conformal time, $\mathrm{d}t\,a^{3}=\mathrm{d}\eta\,a^{4}$, reproducing the scaling assumed in the influence-functional treatment. This reduction should be viewed as the appropriate effective coupling in the mass-dominated long-wavelength regime. For environmental modes that remain deep inside the horizon, $q/a\gg H\gtrsim m_{\sigma}$, the gradient and kinetic terms retained in equation (10) dominate over the mass term, so the subhorizon contribution to decoherence is more generally described by the full operator structure of equation (10) rather than solely by the reduced coupling (11). The coarse-grained estimates below should therefore be interpreted as parametric estimates of decoherence growth, with equation (11) supplying the cleanest explicit coupling in the long-wavelength sector.
For a Gaussian environment state, the noise kernel appearing in equation (5) is [5, 6]
$N(x,x^{\prime})=\lambda_{\mathrm{eff}}^{2}\,\mathrm{Re}\!\left[G^{>}_{\sigma}(x,x^{\prime})^{2}\right],$ (12)
@@
$u_{k}(\eta)=\frac{\sqrt{\pi}}{2}\,\mathrm{e}^{\mathrm{i}(\nu+1/2)\pi/2}\,H(-\eta)^{3/2}\,H^{(1)}_{\nu}(-k\eta),$ (14)
with
$\nu=\sqrt{\frac{9}{4}-\frac{m_{\sigma}^{2}}{H^{2}}},$ (15)
@@
Substituting equations (13)–(14) into equation (12) yields the explicit convolution
$N(\eta,\eta^{\prime};\mathbf{k})=\lambda_{\mathrm{eff}}^{2}\,\mathrm{Re}\!\left[\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\,u_{q}(\eta)\,u^{*}_{q}(\eta^{\prime})\,u_{|\mathbf{k}-\mathbf{q}|}(\eta)\,u^{*}_{|\mathbf{k}-\mathbf{q}|}(\eta^{\prime})\right].$ (17)
@@
$N(\eta,\eta^{\prime};k_{L})\simeq\lambda_{\mathrm{eff}}^{2}\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\mathrm{Re}\!\bigl[\bigl(u_{q}(\eta)\,u^{*}_{q}(\eta^{\prime})\bigr)^{2}\bigr].$ (18)
@@
$N(\eta,\eta^{\prime};k_{L})\propto\lambda_{\mathrm{eff}}^{2}\,H^{4}\,(-\eta)^{2\Delta}(-\eta^{\prime})^{2\Delta}\int\mathrm{d}q\,q^{-4+4\Delta}.$ (19)
This estimate refers to the infrared contribution to the kernel; the horizon-based and EFT-based decoherence estimates below are instead dominated by environmental modes that remain subhorizon over the relevant coarse-graining range. A nonzero mass $m_{\sigma}\neq 0$ softens the infrared behavior, although in the light-field regime the integral still requires a physical infrared cutoff associated with the finite inflationary patch. Writing
$\int_{q_{\min}}\mathrm{d}q\,q^{-4+4\Delta}\propto\frac{q_{\min}^{-3+4\Delta}}{-3+4\Delta},\quad\Delta\neq\tfrac{3}{4},$ (20)
shows explicitly that for any finite $q_{\min}$ the kernel is finite and acquires a regulated dependence on $m_{\sigma}/H$ through $\Delta$. The derivation also shows that the coefficient controlling decoherence is fixed by the spectator-sector Hamiltonian density and therefore has a definite mass dimension; in the minimal model $\lambda_{\mathrm{eff}}\propto m_{\sigma}^{2}$, while in a more general EFT it receives higher-derivative and slow-roll-suppressed corrections.
Special case $\Delta=3/4$.— Equation (20) excludes the value $\Delta=3/4$, corresponding to $\nu=3/4$ and $m_{\sigma}=\tfrac{3\sqrt{3}}{4}H\approx 1.30\,H$. At this mass the integrand $q^{-4+4\Delta}=q^{-1}$ is logarithmically divergent and equation (20) is replaced by
$\int_{q_{\min}}^{q_{\max}}\mathrm{d}q\,q^{-1}=\ln\!\left(\frac{q_{\max}}{q_{\min}}\right),\quad\Delta=\tfrac{3}{4}.$ (21)
The accumulated branch-decoherence functional therefore grows logarithmically in the size of the inflationary patch rather than as a power law, while the overlap factor itself still vanishes superhorizon as $|\mathcal{D}_{k}|\propto z^{3/4}$ — verified numerically to better than $0.03\%$ by comparing $|\mathcal{D}_{k}(z_{1})|/|\mathcal{D}_{k}(z_{2})|$ with $(z_{1}/z_{2})^{3/4}$ at $z_{1}=0.003$, $z_{2}=0.010$, using the general formula (51) with Hankel functions evaluated at $\nu=3/4$. This case lies at the boundary between power-law and logarithmic infrared behavior and is the analogue of the Breitenlohner–Freedman saturating mass for a light field in de Sitter.
III.1 Dissipation kernel and the stochastic inflation limit
The influence action (4) contains both a noise kernel $N$ and a dissipation kernel $D$. While the noise kernel governs decoherence and has been the focus of the preceding analysis, $D$ determines the back-reaction of the environment on the system dynamics and must be computed to obtain the full open-system effective action. We derive $D$ here and show that the resulting Langevin equation for $\zeta_{L}$ recovers Starobinsky’s stochastic inflation [11, 20] in the superhorizon limit, in agreement with the standard derivation [11, 20].
For a Gaussian environment state, the dissipation kernel follows from the imaginary part of the same two-point correlator that gives $N$:
$D(x,x^{\prime})=\lambda_{\mathrm{eff}}^{2}\,\theta(\eta-\eta^{\prime})\,\mathrm{Im}\!\left[G^{>}_{\sigma}(x,x^{\prime})^{2}\right].$ (22)
The $\theta(\eta-\eta^{\prime})$ factor enforces causality: $D$ is a retarded kernel, in contrast to the symmetric noise kernel $N$.
In the WKB approximation for subhorizon environmental modes ($q\gg aH$), the Bunch–Davies mode functions reduce to $u_{q}(\eta)\simeq(2q)^{-1/2}e^{-iq\eta}$, so $(u_{q}u_{q}^{*})^{2}=(4q^{2})^{-1}e^{-2iq\Delta}$ with $\Delta\equiv\eta-\eta^{\prime}$. Substituting into equations (12) and (22) and integrating over the environment ($q>k_{c}$) with an implicit UV convergence factor $e^{-\varepsilon q}$ ($\varepsilon\to 0^{+}$) to regulate the oscillatory tail, one obtains the two kernels in momentum space
$N(\eta,\eta^{\prime};k_{L})$ $=\frac{\lambda_{\mathrm{eff}}^{2}}{64\pi^{2}}\frac{\sin\!\bigl[2k_{c}(\eta-\eta^{\prime})\bigr]}{\eta-\eta^{\prime}},$ (23)
$D(\eta,\eta^{\prime};k_{L})$ $=\frac{\lambda_{\mathrm{eff}}^{2}}{64\pi^{2}}\,\theta(\eta-\eta^{\prime})\,\frac{\cos\!\bigl[2k_{c}(\eta-\eta^{\prime})\bigr]}{\eta-\eta^{\prime}},$ (24)
where $k_{c}=\epsilon\,a(\eta)H$ is the coarse-graining cutoff. The two kernels differ by a $90^{\circ}$ phase shift: $N$ involves $\sin[2k_{c}\Delta]$ while $D$ involves $\cos[2k_{c}\Delta]$, with $\Delta\equiv\eta-\eta^{\prime}$. These are the standard Caldeira–Leggett kernels for an ohmic bath with a sharp lower cutoff [6, 21].
Fluctuation-dissipation relation.— In the long-wavelength limit $|\omega|\ll k_{c}$, the Fourier transforms of equations (23)–(24) are consistent with the de Sitter fluctuation-dissipation relation
$\tilde{D}(\omega)$ $=-\frac{\mathrm{i}\omega}{2T_{\mathrm{GH}}}\,\tilde{N}(\omega),$
$T_{\mathrm{GH}}$ $=\frac{H}{2\pi},$ (25)
where $T_{\mathrm{GH}}$ is the Gibbons–Hawking temperature of de Sitter space [24]. This KMS-type relation is a general property of quantum fields in de Sitter and holds for the exact Bunch–Davies two-point function; the WKB kernels above are consistent with the interpretation that the open-system dynamics is tied to the horizon thermodynamics at $T_{\mathrm{GH}}=H/(2\pi)$.
Stochastic equation of motion.— The influence action with both $N$ and $D$ generates a Langevin equation for the long-wavelength curvature perturbation,
$\mathcal{P}[\zeta_{L}](\eta)+\Gamma\,\partial_{\eta}\zeta_{L}(\eta)=\xi(\eta),$ (26)
where $\mathcal{P}$ is the free equation-of-motion operator for $\zeta_{L}$ in de Sitter (which reduces to $\epsilon H^{2}\zeta_{L}$ at leading order in slow roll on superhorizon scales), $\xi(\eta)$ is a Gaussian stochastic noise satisfying
$\langle\xi(\eta)\rangle$ $=0,$
$\langle\xi(\eta)\,\xi(\eta^{\prime})\rangle$ $=N(\eta,\eta^{\prime};k_{L}),$ (27)
and the friction coefficient is estimated in the Markovian approximation. The integral $\int_{0}^{\infty}D(\Delta)\mathrm{d}\Delta$ is oscillatory and requires UV regularization; with a physical UV cutoff $q_{\max}\gg k_{c}$, the leading contribution is
$\Gamma\simeq\frac{\lambda_{\mathrm{eff}}^{2}}{128\pi^{2}k_{c}}=\frac{\hat{\lambda}^{2}H}{128\pi^{2}\epsilon},$ (28)
where the $k_{c}$ in the denominator arises because the dominant contribution to the Markovian integral comes from the scale $\Delta\sim 1/(2k_{c})$ set by the coarse-graining cutoff. The numerical prefactor in equation (28) is regulator-dependent at order unity; the robust content is the scaling $\Gamma\propto\lambda_{\mathrm{eff}}^{2}/k_{c}$. Since $\Gamma/H=\hat{\lambda}^{2}/(128\pi^{2}\epsilon)\ll 1$ for $\hat{\lambda}\ll 1$ and $\epsilon\sim\mathcal{O}(1)$, the friction term is negligible on superhorizon scales and $\zeta_{L}$ is effectively frozen. In the superhorizon limit $\mathcal{P}[\zeta_{L}]\to 0$ and, using the local (white-noise) approximation $N(\eta,\eta^{\prime})\approx(H^{2}/4\pi^{2})\,a^{-4}\,\delta(\eta-\eta^{\prime})$, the Langevin equation (26) reduces to
$\mathrm{d}\zeta_{L}\simeq\frac{H}{2\pi}\,\mathrm{d}W(\mathcal{N}),$ (29)
where $\mathrm{d}W$ is a Wiener increment per e-fold. This is Starobinsky’s stochastic inflation equation [11, 20], recovered here as the superhorizon limit of the Langevin equation (26), in agreement with the standard derivation [11, 20]. The influence functional therefore provides a systematic open-system framework connecting the decoherence calculation above to the classical stochastic description of inflationary perturbations.
III.2 Coarse graining and decoherence rates